# Imitation from Diverse Behaviors: Wasserstein Quality Diversity Imitation Learning with Single-Step Archive Exploration

### v1

@@ Abstract
Learning diverse and high-performance behaviors from a limited set of demonstrations is a grand challenge. Traditional imitation learning methods usually fail in this task because most of them are designed to learn one specific behavior even with multiple demonstrations. Therefore, novel techniques for quality diversity imitation learning are needed to solve the above challenge. This work introduces Wasserstein Quality Diversity Imitation Learning (WQDIL), which 1) improves the stability of imitation learning in the quality diversity setting with latent adversarial training based on a Wasserstein Auto-Encoder (WAE), and 2) mitigates a behavior-overfitting issue using a measure-conditioned reward function with a single-step archive exploration bonus. Empirically, our method significantly outperforms state-of-the-art IL methods, achieving near-expert or beyond-expert QD performance on the challenging continuous control tasks derived from MuJoCo environments.
@@
Keywords Imitation Learning, Quality Diversity, Wasserstein Auto-Encoder, Wasserstein Adversarial Training, Single-Step Archive Exploration
1 Introduction
@@
Imitation Learning (IL) aims to mimic an expert’s behavior by learning from the demonstrations. IL has achieved great success in many real-world applications, such as robotics [51], autonomous driving [6], and drone control [37]. However, most of the traditional imitation learning methods were designed to learn one specific behavior, even given multiple demonstrations.
@@
Learning diverse and high-quality policies is the ultimate goal of many real-world applications like robotic locomotion tasks [4]. Recent literature has demonstrated that quality diversity reinforcement learning (QDRL) [4, 44] is a promising direction for achieving this goal. Prior methods that combine Differentiable Quality Diversity (DQD) [14] with off-policy RL achieved diverse and relatively high-performance policies. However, the performance gap between standard RL and QDRL still exists [4]. More recently, Batra et al. [4] mitigate this gap by leveraging the on-policy RL method PPO [39] with DQD. PPO estimates the gradients of diversity and performance objectives from the online collected data. Then, the estimated gradients are used by DQD methods like CMA-MAEGA [15] that maintain a single search point and move through the behavior space by filling new regions. The synergy between PPO and DQD results in a state-of-the-art QDRL method, Proximal Policy Gradient Arborescence (PPGA) [4], that achieves the ultimate goal of robotic locomotion. Generally, this kind of QDRL method finds a diverse archive of high-performing locomotion behaviors for an agent by combining PPO gradient approximations with Differentiable Quality Diversity algorithms.
@@
However, the success of QDRL heavily relies on high-quality reward functions, which can be intractable in practice. Quality Diversity Imitation Learning (QDIL) offers a more flexible strategy for learning diverse and high-quality policies from demonstrations with diverse behaviors. In the literature, adversarial IL methods such as GAIL [23] have achieved great success in learning specific behaviors for robotic locomotion tasks. Therefore, a naive solution for QDIL is to apply adversarial IL to estimate rewards from the given demonstrations and then leverage the estimated rewards for learning diverse and high-quality policies. Unfortunately, adversarial IL methods suffer from the training instability issue, which usually results in worse-than-demonstrator performance [23]. Moreover, when the demonstrations only contain a few behaviors, the rewards learned by adversarial IL techniques will be behavior-overfitted and unable to guide the agent to learn more diverse behaviors beyond the demonstrations.
@@
Figure 1: Illustrations of the two issues of the Adversarial QDIL (i.e., training instability and behavior-overfitted reward) and their corresponding solutions (i.e. WQDIL and Single-Step Archive Exploration). $\delta(s)$ means the Markovian Measure Proxy of state $s$, a.k.a. the single-step measure.
@@
The training instability issue and the behavior-overfitted reward issue heavily limit the adversarial QDIL to learning diverse and high-quality policies with limited demonstrations. In this work, we propose two synergic strategies to overcome these two challenging issues. The first strategy aims to stabilize the training of the reward model, and the second strategy focuses on encouraging the agent to do behavior space exploration. To develop the first strategy, we propose to stabilize the reward learning by applying Wasserstein adversarial training within the latent space of the Wasserstein Auto-Encoder (WAE) [46]. Similar to VAE [26], WAE keeps the good properties of stable training and a nice latent manifold structure while generating higher-quality images than GAN [46]. Therefore, we propose to apply WAE to enable a more stable training of reward model in adversarial QDIL. In addition, we propose latent Wasserstein adversarial training to further improve the consistency of the reward training stability. For the second strategy, we introduce a bonus that enables the agent to collect data with more diverse behaviors via single-step archive exploration. We call the resulting method Wasserstein Quality Diversity Imitation Learning (WQDIL) with Single-Step Archive Exploration (SSAE). Figure 1 illustrates the two issues of the Adversarial QDIL (i.e., training instability and behavior-overfitted reward) and the corresponding solutions (i.e. WQDIL and Single-Step Archive Exploration). The synergy between WQDIL and SSAE adherently address the two issues of adversarial QDIL, resulting in diverse and high-quality policies when learning with limited demonstrations. Figure 2 illustrates diverse behaviors induced by our learned policies.
Figure 2: Illustration of diverse behaviors learned by our Quality Diversity Imitation Learning framework on Humanoid and Walker2d, where each column represents one behavior. The “left” and “‘right” means the proportion of time the left leg or right leg contacting the ground.
@@
  • Empirically, we demonstrated that our method significantly outperforms state-of-the-art imitation learning methods in learning diverse and high-quality policies from limited demonstrations.
2 Background and Related Work
@@
2.1 Reinforcement Learning and Imitation Learning
@@
Imitation learning (IL) trains an agent to mimic expert behaviors from demonstrations [50]. Behavior Cloning (BC) uses supervised learning to imitate expert behavior but suffers from severe error accumulation [36]. Inverse Reinforcement Learning (IRL) seeks to recover a reward function from the demonstrations, and use reinforcement learning (RL) to train a policy that best mimics the expert behaviors [1]. Early IRL methods estimate rewards in the principle of maximum entropy [52, 48, 13]. Adversarial Inverse Reinforcement Learning (AIRL) [18] learns a robust reward function by training the discriminator via logistic regression to classify expert data from policy data.
@@
Adversarial IL methods treat IRL as a distribution-matching problem. Generative Adversarial Imitation Learning (GAIL) [23] trains a discriminator to differentiate between the state-action distribution of the demonstrations and the state-action distribution induced by the agent’s policy, and output a reward to guide policy improvement. More recently, Primal Wasserstein Imitation Learning (PWIL) [12] introduces an offline reward function based on an upper bound of the Wasserstein distance between the expert and agent’s state-action distributions, avoiding the instability of adversarial IL methods. This paper includes max-entropy IRL (MaxEntIRL), the classic adversarial IL (i.e., GAIL) and the state-of-the-art PWIL as our baselines in QDIL with limited demonstrations.
2.2 Quality Diversity Reinforcement Learning
@@
Quality Diversity Optimization. Distinct from traditional optimization which aims to find a single solution to maximize the objective, Quality Diversity (QD) optimization aims to find a set of high-quality and diverse solutions in an $n$-dimensional continuous space $\mathbb{R}^{n}$. Given an objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and $k$-dimensional measure function $m:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$, the goal is to find solutions $\theta\in\mathbb{R}^{n}$ for each local region in the behavior space $B=m(\mathbb{R}^{n})$. QD algorithms discretize $B$ into $M$ cells, forming an archive $\mathcal{A}$. Formally, the objective is to find a set of solutions $\{\theta_{i}\}_{i=1}^{M}$ which maximises $f(\theta_{i})$ for each $i=1,\dots,M$. Each solution $\theta_{i}$ corresponds to a cell in $\mathcal{A}$ via its measure $m(\theta_{i})$, forming an archive of high-quality and diverse solutions [10, 34].
@@
Previous Quality Diversity optimization methods integrate Evolution Strategies (ES) with MAP-Elites [31], such as Covariance Matrix Adaptation MAP-Elites (CMA-ME) [16]. CMA-ME uses CMA-ES [21] as an ES algorithm generating new solutions inserted into the archive, and uses MAP-Elites to retain the highest-performing solution in each cell. CMA-ES adapts its sampling distribution based on archive improvements from offspring solutions. However, traditional ES faces low sample efficiency, especially for high-dimensional parameters such as neural networks. Differentiable Quality Diversity (DQD) enhances exploration and optimization by leveraging gradient information from both the objective and measure functions. The state-of-the-art DQD algorithm, CMA-MAEGA [15], utilizes CMA-ES to maintain a distribution over coefficient sets $c_{i}$, rather than directly generating solutions $\theta$. The objective function is $g(\theta)=|c_{0}|f(\theta)+\sum_{j=1}^{k}c_{j}m_{j}(\theta)$, where coefficients $c_{i}$ weight the gradients of the objective $f$ and measures $m_{j}$. CMA-MAEGA maintains a search policy and samples multiple coefficient sets after computing gradients of $f$ and $m_{i}$ to generate branched offspring solutions $\pi_{\theta_{1}},\dots,\pi_{\theta_{\lambda}}$. The archive improvement from each offspring guides the update of the coefficient distribution, which is then used to refine the search policy.
@@
Built on the QD optimization paradigm, the Quality Diversity Reinforcement Learning (QDRL) problem can be viewed as maximizing $f(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{k=0}^{T-1}\gamma^{k}r(s_{k},a_{k})\right]$ with respect to diverse $\theta$ in a policy archive defined by measure $m$ [11]. In QDRL, both the objective and measure are non-differentiable, requiring approximations by DQD approaches. The state-of-the-art QDRL algorithm, Proximal Policy Gradient Arborescence (PPGA), employs a vectorized PPO (VPPO) architecture to approximate the gradients of the objective and measure functions [5]. PPGA introduced the Markovian Measure Proxy (MMP), a surrogate measure function that correlates strongly with the original measure and allows gradient approximation via policy gradient by treating it as a reward function. Specifically, MMP decomposes trajectory-based measures $m$ into individual steps by computing:
@@
2.3 Quality Diversity Imitation Learning
Definition 1 (Quality-Diversity Imitation Learning (QDIL))
@@ 2.3 Quality Diversity Imitation Learning
In this paper, we focus on Adversarial QDIL, in which we apply adversarial IL, based on GAIL [19] in particular, to estimate rewards from the given demonstrations $\mathcal{D}$. This adversarial training scheme considers the policy as an adversary and an additional neural network, the disciminator $D$, which discriminates between samples of the policy and samples of the demonstrations. The resulting reward model is based on the discriminator, and indicates the likelihood of corresponding to demonstration behaviours. This reward model then is used for the objective of QDIL.
2.4 Improving Stability of Adversarial IL
@@
The methodology of adversarial IL comes from Generative Adversarial Networks (GAN) [19]. Although adversarial IL methods like GAIL have achieved great success in low-dimensional environments, they also inherit the training instability issue of GAN that limits its generalization to more diverse tasks [28].
@@
Wasserstein GAN (WGAN) [2] was proposed to stabilize the training of GAN by minimizing the Wasserstein distance instead of the Jensen-Shannon divergence between true data distribution and generated data distribution. However, the original WGAN with weight clipping can lead to undesired behaviors [20]. Gradient penalty (GP) was later proposed to address this issue, and has become a widely-used strategy for improving the stability of GAN and WGAN [38, 33, 20, 30, 43, 35]. We have applied these stabilization methods in adversarial QDIL, and conduct preliminary investigation experiments on continuous control tasks derived from MuJoCo. However, we only observed negligible effects of these methods on stabilizing adversarial QDIL.
@@
Wasserstein Auto-Encoder (WAE) [46] minimizes the optimal transport cost $W_{c}(P_{X},P_{G})$ based on the novel auto-encoder formulation. In the resulting optimization problem, the decoder tries to accurately reconstruct the encoded training examples as measured by the cost function $f_{c}$. Similar to VAE [26], the encoder tries to simultaneously achieve two conflicting goals: it tries to match the encoded distribution of training examples $Q_{Z}:=\mathbb{E}_{P_{X}}[Q(Z|X)]$ to the prior $P_{Z}$ as measured by any specified divergence $D_{z}(Q_{Z},P_{Z})$, while making sure that the latent codes provided to the decoder are informative enough to reconstruct the encoded training examples. Using the squared cost $f_{c}(x,y)=\|x-y\|_{2}^{2}$, WAE keeps the good properties of VAEs (stable training, and a nice latent manifold structure) while generating better-quality images than GAN [46]. This observation inspired us to apply WAE in improving the stability of adversarial QDIL. The resulting framework is called Wasserstein Quality Diversity Imitation Learning (WQDIL).
2.5 Encouraging exploration in RL and IL

### v2

@@ Abstract.
Learning diverse and high-performance behaviors from a limited set of demonstrations is a grand challenge. Traditional imitation learning methods usually fail in this task because most of them are designed to learn one specific behavior even with multiple demonstrations. Therefore, novel techniques for quality diversity imitation learning, which bridges the quality diversity optimization and imitation learning methods, are needed to solve the above challenge. This work introduces Wasserstein Quality Diversity Imitation Learning (WQDIL), which 1) improves the stability of imitation learning in the quality diversity setting with latent adversarial training based on a Wasserstein Auto-Encoder (WAE), and 2) mitigates a behavior-overfitting issue using a measure-conditioned reward function with a single-step archive exploration bonus. Empirically, our method significantly outperforms state-of-the-art IL methods, achieving near-expert or beyond-expert QD performance on the challenging continuous control tasks derived from MuJoCo environments.
Key words and phrases:
@@
Imitation Learning, Quality Diversity, Wasserstein Auto-Encoder, Wasserstein Adversarial Training, Single-Step Archive Exploration
1. Introduction
@@
Imitation Learning (IL) aims to mimic an expert’s behavior by learning from the demonstrations. IL has achieved great success in many real-world applications, such as robotics (Zhu et al., 2018), autonomous driving (Bojarski, 2016), and drone control (Ross et al., 2013). However, most of the traditional imitation learning methods were designed to learn one specific behavior, even given multiple demonstrations.
@@
Learning diverse and high-quality policies is the ultimate goal of many real-world applications like robotic locomotion tasks (Batra et al., 2023a). Recent literature has demonstrated that quality diversity reinforcement learning (QDRL) (Batra et al., 2023a; Tjanaka et al., 2022) is a promising direction for achieving this goal. Prior methods that combine Differentiable Quality Diversity (DQD) (Fontaine and Nikolaidis, 2021) with off-policy RL achieved diverse and relatively high-performance policies. However, the performance gap between standard RL and QDRL still exists (Batra et al., 2023a). More recently, Batra et al. (Batra et al., 2023a) mitigate this gap by leveraging the on-policy RL method PPO (Schulman et al., 2017) with DQD. PPO estimates the gradients of diversity and performance objectives from the online collected data. Then, the estimated gradients are used by DQD methods like CMA-MAEGA (Fontaine and Nikolaidis, 2023) that maintain a single search point and move through the behavior space by filling new regions. The synergy between PPO and DQD results in a state-of-the-art QDRL method, Proximal Policy Gradient Arborescence (PPGA) (Batra et al., 2023a), that achieves the ultimate goal of robotic locomotion. Generally, this kind of QDRL method finds a diverse archive of high-performing locomotion behaviors for an agent by combining PPO gradient approximations with Differentiable Quality Diversity algorithms.
@@
However, the success of QDRL heavily relies on high-quality reward functions, which can be intractable in practice. Quality Diversity Imitation Learning (QDIL) offers a more flexible strategy for learning diverse and high-quality policies from demonstrations with diverse behaviors. In the literature, adversarial IL methods such as GAIL (Ho and Ermon, 2016) have achieved great success in learning specific behaviors for robotic locomotion tasks. Therefore, a naive solution for QDIL is to apply adversarial IL to estimate rewards from the given demonstrations and then leverage the estimated rewards for learning diverse and high-quality policies. Unfortunately, adversarial IL methods suffer from the training instability issue, which usually results in worse-than-demonstrator performance (Ho and Ermon, 2016). Moreover, when the demonstrations only contain a few behaviors, the rewards learned by adversarial IL techniques will be behavior-overfitted and unable to guide the agent to learn more diverse behaviors beyond the demonstrations.
@@
Figure 1. Illustration of the two issues of the Adversarial QDIL (i.e., training instability and behavior-overfitted reward) and their corresponding solutions (i.e. WQDIL and Single-Step Archive Exploration). $\delta(s)$ means the Markovian Measure Proxy of state $s$, a.k.a. the single-step measure.
@@
The training instability issue and the behavior-overfitted reward issue heavily limit the adversarial QDIL to learning diverse and high-quality policies with limited demonstrations. In this work, we propose two synergic strategies to overcome these two challenging issues. The first strategy aims to stabilize the training of the reward model, and the second strategy focuses on encouraging the agent to do behavior space exploration. To develop the first strategy, we propose to stabilize the reward learning by applying Wasserstein adversarial training within the latent space of the Wasserstein Auto-Encoder (WAE) (Tolstikhin et al., 2018). Similar to VAE Kingma and Welling (2014), WAE keeps the good properties of stable training and a nice latent manifold structure while generating higher-quality images than GAN Tolstikhin et al. (2018). Therefore, we propose to apply WAE to enable a more stable training of reward model in adversarial QDIL. In addition, we propose latent Wasserstein adversarial training to further improve the consistency of the reward training stability. For the second strategy, we introduce a bonus that enables the agent to collect data with more diverse behaviors via single-step archive exploration. We call the resulting method Wasserstein Quality Diversity Imitation Learning (WQDIL) with Single-Step Archive Exploration (SSAE). Figure 1 illustrates the two issues of the Adversarial QDIL (i.e., training instability and behavior-overfitted reward) and the corresponding solutions (i.e. WQDIL and Single-Step Archive Exploration). The synergy between WQDIL and SSAE adherently address the two issues of adversarial QDIL, resulting in diverse and high-quality policies when learning with limited demonstrations.
@@
Figure 2. Illustration of diverse behaviors learned by our Quality Diversity Imitation Learning framework on Humanoid and Walker2d, where each column represents one behavior. The “left” and “‘right” means the proportion of time the left leg or right leg contacting the ground.
2. Background and Related Work
@@
2.1. Reinforcement Learning and Imitation Learning
@@
Imitation learning (IL) trains an agent to mimic expert behaviors from demonstrations (Zare et al., 2024). Behavior Cloning (BC) uses supervised learning to imitate expert behavior but suffers from severe error accumulation (Ross et al., 2011). Inverse Reinforcement Learning (IRL) seeks to recover a reward function from the demonstrations, and use reinforcement learning (RL) to train a policy that best mimics the expert behaviors (Abbeel and Ng, 2004). Early IRL methods estimate rewards in the principle of maximum entropy (Ziebart et al., 2008; Wulfmeier et al., 2015; Finn et al., 2016). Adversarial Inverse Reinforcement Learning (AIRL) (Fu et al., 2017) learns a robust reward function by training the discriminator via logistic regression to classify expert data from policy data.
@@
Adversarial IL methods treat IRL as a distribution-matching problem. Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016) trains a discriminator to differentiate between the state-action distribution of the demonstrations and the state-action distribution induced by the agent’s policy, and output a reward to guide policy improvement. More recently, Primal Wasserstein Imitation Learning (PWIL) (Dadashi et al., 2021) introduces an offline reward function based on an upper bound of the Wasserstein distance between the expert and agent’s state-action distributions, avoiding the instability of adversarial IL methods. This paper includes max-entropy IRL (MaxEntIRL), the classic adversarial IL (i.e., GAIL) and the state-of-the-art PWIL as our baselines in QDIL with limited demonstrations.
2.2. Quality Diversity Reinforcement Learning
@@
Quality Diversity Optimization. Distinct from traditional optimization which aims to find a single solution to maximize the objective, Quality Diversity (QD) optimization aims to find a set of high-quality and diverse solutions in an $n$-dimensional continuous space $\mathbb{R}^{n}$. Given an objective function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and $k$-dimensional measure function $m:\mathbb{R}^{n}\rightarrow\mathbb{R}^{k}$, the goal is to find solutions $\theta\in\mathbb{R}^{n}$ for each local region in the behavior space $B=m(\mathbb{R}^{n})$. QD algorithms discretize $B$ into $M$ cells, forming an archive $\mathcal{A}$. Formally, the objective is to find a set of solutions $\{\theta_{i}\}_{i=1}^{M}$ which maximises $f(\theta_{i})$ for each $i=1,\dots,M$. Each solution $\theta_{i}$ corresponds to a cell in $\mathcal{A}$ via its measure $m(\theta_{i})$, forming an archive of high-quality and diverse solutions (Chatzilygeroudis et al., 2021; Pugh et al., 2016).
@@
Previous Quality Diversity optimization methods integrate Evolution Strategies (ES) with MAP-Elites (Mouret and Clune, 2015), such as Covariance Matrix Adaptation MAP-Elites (CMA-ME) (Fontaine et al., 2020). CMA-ME uses CMA-ES (Hansen and Ostermeier, 2001) as an ES algorithm generating new solutions inserted into the archive, and uses MAP-Elites to retain the highest-performing solution in each cell. CMA-ES adapts its sampling distribution based on archive improvements from offspring solutions. However, traditional ES faces low sample efficiency, especially for high-dimensional parameters such as neural networks. Differentiable Quality Diversity (DQD) enhances exploration and optimization by leveraging gradient information from both the objective and measure functions. The state-of-the-art DQD algorithm, CMA-MAEGA (Fontaine and Nikolaidis, 2023), utilizes CMA-ES to maintain a distribution over coefficient sets $c_{i}$, rather than directly generating solutions $\theta$. The objective function is $g(\theta)=|c_{0}|f(\theta)+\sum_{j=1}^{k}c_{j}m_{j}(\theta)$, where coefficients $c_{i}$ weight the gradients of the objective $f$ and measures $m_{j}$. CMA-MAEGA maintains a search policy and samples multiple coefficient sets after computing gradients of $f$ and $m_{i}$ to generate branched offspring solutions $\pi_{\theta_{1}},\dots,\pi_{\theta_{\lambda}}$. The archive improvement from each offspring guides the update of the coefficient distribution, which is then used to refine the search policy.
@@
Built on the QD optimization paradigm, the Quality Diversity Reinforcement Learning (QDRL) problem can be viewed as maximizing $f(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{k=0}^{T-1}\gamma^{k}r(s_{k},a_{k})\right]$ with respect to diverse $\theta$ in a policy archive defined by measure $m$ (Cideron et al., 2020). In QDRL, both the objective and measure are non-differentiable, requiring approximations by DQD approaches. The state-of-the-art QDRL algorithm, Proximal Policy Gradient Arborescence (PPGA), employs a vectorized PPO (VPPO) architecture to approximate the gradients of the objective and measure functions (Batra et al., 2023b). PPGA introduced the Markovian Measure Proxy (MMP), a surrogate measure function that correlates strongly with the original measure and allows gradient approximation via policy gradient by treating it as a reward function. Specifically, MMP decomposes trajectory-based measures $m$ into individual steps by computing:
@@
2.3. Quality Diversity Imitation Learning
Definition 1 (Quality-Diversity Imitation Learning (QDIL)).
@@ 2.3. Quality Diversity Imitation Learning
In this paper, we focus on Adversarial QDIL, in which we apply adversarial IL, based on GAIL Goodfellow et al. (2014) in particular, to estimate rewards from the given demonstrations $\mathcal{D}$. This adversarial training scheme considers the policy as an adversary and an additional neural network, the disciminator $D$, which discriminates between samples of the policy and samples of the demonstrations. The resulting reward model is based on the discriminator, and indicates the likelihood of the parameters based on the demonstration behaviours. This reward model is then used for the objective of QDIL.
2.4. Improving Stability of Adversarial IL
@@
The methodology of adversarial IL comes from Generative Adversarial Networks (GAN) (Goodfellow et al., 2014). Although adversarial IL methods like GAIL have achieved great success in low-dimensional environments, they also inherit the training instability issue of GAN that limits its generalization to more diverse tasks (Luo et al., 2024). This problem will be exacerbated in the QDIL setting due to the variety of policies being trained.
@@
Wasserstein GAN (WGAN) (Arjovsky et al., 2017) was proposed to stabilize the training of GANs by minimizing the Wasserstein distance instead of the Jensen-Shannon divergence between true data distribution and generated data distribution. However, the original WGAN with weight clipping can lead to undesired behaviors (Gulrajani et al., 2017). Gradient penalty (GP) was later proposed to address this issue, and has become a widely-used strategy for improving the stability of GAN and WGAN (Roth et al., 2017; Petzka et al., 2017; Gulrajani et al., 2017; Mescheder et al., 2018; Thanh-Tung et al., 2019; Qi, 2020). We have applied these stabilization methods in adversarial QDIL, and conduct preliminary investigation experiments on continuous control tasks derived from MuJoCo. However, we only observed negligible effects of these methods on stabilizing adversarial QDIL.
@@
Wasserstein Auto-Encoder (WAE) (Tolstikhin et al., 2018) minimizes the optimal transport cost $W_{c}(P_{X},P_{G})$ based on the novel auto-encoder formulation. In the resulting optimization problem, the decoder tries to accurately reconstruct the encoded training examples as measured by the cost function $f_{c}$. Similar to VAE (Kingma and Welling, 2014), the encoder tries to simultaneously achieve two conflicting goals: it tries to match the encoded distribution of training examples $Q_{Z}:=\mathbb{E}_{P_{X}}[Q(Z|X)]$ to the prior $P_{Z}$ as measured by any specified divergence $D_{z}(Q_{Z},P_{Z})$, while making sure that the latent codes provided to the decoder are informative enough to reconstruct the encoded training examples. Using the squared cost $f_{c}(x,y)=\|x-y\|_{2}^{2}$, WAE keeps the good properties of VAEs (stable training and a nice latent manifold structure) while generating better-quality images than GAN (Tolstikhin et al., 2018). This observation inspired us to apply WAE to improve the stability of adversarial QDIL. The resulting framework is called Wasserstein Quality Diversity Imitation Learning (WQDIL).
2.5. Encouraging exploration in RL and IL